OpenStudy (anonymous):

Four of Ethan's family members went on a 4 and two-thirds hour road trip. If they all shared the drive time equally, how many hours did each family member drive? A)SIX-SEVENTHS B)1 AND TWO-THIRDS C)ELEVEN-TWELFTHS D)1 AND ONE-SIXTH SOMEONE HELP ME PLZ????

OpenStudy (texaschic101):

(4 2/3) / 4 = (14/3) / 4 = 14/3 * 1/4 = 14/12 = 1 1/6

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

NAWW THANX

OpenStudy (texaschic101):

no problem...happy to help :)

OpenStudy (anonymous):

yah i feel like some one is tryna you know stock me or something not saying u but i looked at my message nd someone is tryna you know sauce me nd i dont appreciate dat